### v2

# Vertex Ordering Characterizations of Interval $r$-Graphs

Abstract
An $r$-partite graph is an interval $r$-graph if corresponding to each vertex we can assign an interval of the real line such that two vertices $u$ and $v$ of different partite sets are adjacent if and only if their corresponding intervals intersect. In this paper, we provide two vertex-ordering characterizations of interval $r$-graphs, utilizing generalized interval orderings and $r$-interval orderings, and identify forbidden patterns for interval $r$-graphs in terms of specific orderings of their vertices.
Keywords: interval $r$-graphs, vertex ordering, generalized interval ordering, $r$-interval ordering , forbidden patterns
1 Introduction
A graph $G=(V,E)$ is called an interval graph, if there exists a family $\mathcal{I}=\{I_{v}:v\in V\}$, of intervals of the real line such that, for $a,b\in V$ the vertices $a$ and $b$ are adjacent in $G$ if and only if $I_{a}\cap I_{b}\neq\emptyset$. A bigraph $B=(X,Y,E)$ is said to be an interval bigraph if there exists a family $\mathcal{I}=\{I_{v}:v\in X\cup Y\}$ of intervals of the real line such that for all $x\in X$ and $y\in Y$ the vertices $x$ and $y$ are adjacent in $B$ if and only if $I_{x}\cap I_{y}\neq\emptyset$. Interval bigraphs were introduced in [5] and have been studied in [2, 3, 6, 7]. Interval digraphs were introduced by Sen et al. [8]. It was shown in [3] that the notions of interval digraphs and interval bigraphs are equivalent. Interval bigraphs or interval digraphs have become of interest in such new areas as graph homomorphism, e.g [4]. The concept of interval bigraph was generalized by Brown [1]. A graph $G=(V,E)$ is called interval $r$-graph if the vertex set $V$ can be partitioned into $r$ stable sets say $V_{1}$, $V_{2}$,…, $V_{r}$ and corresponding to each vertex we can assign an interval such that two vertices $u$ and $v$ are adjacent in $G$ if and only if the corresponding intervals $I_{u}$ and $I_{v}$ have non-empty intersection and $u$ and $v$ belong to different partite sets of vertices.
In this paper, we study different types of vertex orderings of interval $r$-graphs and, with the help of these orderings, we characterize the class of interval $r$-graphs. Finally, we provide forbidden patterns for interval $r$-graphs with respect to a particular vertex ordering.
2 Main Result
In this section, we introduce several types of vertex orderings for $r$-partite graphs and using these orderings we characterize interval $r$-graphs. Based on these orderings, we provide characterizations and identify forbidden patterns for interval $r$-graphs.
We begin with the definition of generalized interval ordering of the vertices of an $r$-partite graph.
Definition 1.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph of order $n$. Arrange the vertices of $G$ linearly and label them $v_{1},v_{2},\ldots,v_{n}$ such that the vertex $v_{i}$ is placed on the $i$-th positive integer of the real line. Let $X=\bigcup_{i=1}^{r}X_{i}$ denote the vertex set of $G$. The set $X$ is said to admit a generalized interval ordering if the following condition is satisfied:
  • For any edge $v_{i}v_{j}\in E$ with $i>j$, where $v_{i}\in X_{\alpha}$, $v_{j}\in X_{\beta}$, and $\alpha\neq\beta$, it holds that $v_{\ell}v_{j}\in E\quad\text{for all }v_{\ell}\notin X_{\beta}\text{ with }\ell\in\{j+1,j+2,\ldots,i-1\}.$
We shall now characterize interval $r$-graphs using this ordering.
Theorem 1.
An $r$-partite graph $G=(X_{1},X_{2},...X_{r},E)$ is a interval $r$-graph if and only if the vertex set $X=\bigcup_{i=1}^{r}X_{i}$ of $G$ admits a generalized interval ordering.
Proof.
Necessity: Let $G=(X_{1},X_{2},\dots,X_{r},E)$ be an interval $r$-graph. Then there exists a class $\mathcal{I}=\{I_{v}:v\in\bigcup_{i=1}^{r}X_{i}\}$ of intervals corresponding to each vertex $v$ of $G$, such that $uv\in E$ if and only if $I_{u}\cap I_{v}\neq\emptyset$, where $u$ and $v$ belong to different partite sets. Without loss of generality, we assume that all intervals have distinct endpoints.
Now, order the vertices of $G$ as $v_{1},v_{2},\dots,v_{n}$ according to the increasing order of the left endpoints of their corresponding intervals, where $n$ denotes the order of the $r$-partite graph $G$. We will show that this ordering is a generalized interval ordering for the $r$-partite graph $G$.
Suppose $v_{i}v_{j}\in E$, where $i>j$ (and $v_{i}\in X_{\alpha}$, $v_{j}\in X_{\beta}$, $\alpha\neq\beta$). According to our ordering, the left endpoint of $v_{i}$ must lie within the interval $I_{v_{j}}$ corresponding to $v_{j}$. From Figure 1, it follows that the left endpoints of the intervals $I_{v_{j+1}},I_{v_{j+2}},I_{v_{j+3}},\dots,I_{v_{i-1}}$ are all contained in $I_{v_{j}}$. Thus $v_{l}v_{j}\in E$ if $v_{l}\notin X_{\beta}$ for $l\in\{j+1,\dots,i-1\}$. Hence, by Definition 1, we conclude that the ordering $v_{1},v_{2},\dots,v_{n}$ of the vertices of $G$ is a generalized interval ordering.
$v_{1}$$v_{2}$$v_{j}$$v_{j+1}$$v_{j+2}$$v_{i-1}$$v_{i}$$v_{i+1}$$v_{n-1}$$v_{n}$$I_{v_{i}}$$I_{v_{j}}$
Figure 1: Left end point of $I_{v_{i}}$ is contained in $I_{v_{j}}$.
Sufficiency: Let $G=(X_{1},X_{2},\dots,X_{r},E)$ be an $r$-partite graph and the ordering $v_{1},v_{2},\dots,v_{n}$ of its vertices is a generalized interval ordering.
Now, let $m_{i}=\max\left(\{\,j:v_{i}v_{j}\in E\,\}\cup\{i\}\right)$. Construct an interval $I^{\prime}_{v_{i}}$ for each vertex $v_{i}$ $(1\leq i\leq n)$ as $I^{\prime}_{v_{i}}=\left[\,i,m_{i}+\frac{1}{2}\,\right]$. First we shall show that the collection $\mathcal{I}^{\prime}=\{\,I^{\prime}_{v_{i}}:1\leq i\leq n\,\}$ of intervals satisfies the necessary conditions for $G$ to be an interval $r$-graph.
Let $v_{i}$ and $v_{j}$ be two vertices belonging to different partite sets of $G$ such that $v_{i}v_{j}\in E$ and $i>j$. According to our construction, $I^{\prime}_{v_{i}}=\left[\,i,m_{i}+\frac{1}{2}\,\right]\quad\text{and}\quad I^{\prime}_{v_{j}}=\left[\,j,m_{j}+\frac{1}{2}\,\right]$. Since $v_{i}v_{j}\in E$ and $i>j$, we have $m_{j}\geq i$, which implies $m_{j}+\frac{1}{2}\geq i$. Therefore, the interval $I^{\prime}_{v_{j}}$ contains $i$, and hence $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{j}}\neq\emptyset$.
Conversely, let $v_{i}$ and $v_{j}$ be two vertices of $G$ from different partite sets such that $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{j}}\neq\emptyset$ and $i>j$. Then $i\in I^{\prime}_{v_{j}}$, which implies $m_{j}\geq i$. Thus, $j<i\leq m_{j}$, and since $v_{j}v_{m_{j}}\in E$, by the definition of a generalized interval ordering we have $v_{i}v_{j}\in E$. Hence, $v_{i}v_{j}\in E\quad\Longleftrightarrow\quad I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{j}}\neq\emptyset$. Therefore, $G$ is an interval $r$-graph.
Before presenting another vertex-ordering characterization, we define the concept of almost consecutive ones in the rows and columns of the adjacency matrix of $r$-partite graphs.
Definition 2.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph and let A be the adjacency matrix of $G$. A row (say, the $i$-th row) is said to have almost consecutive ones if, between any two ones in that row, whenever a zero appears at the position $(v_{i},v_{k})$, the vertices $v_{i}$ and $v_{k}$ belong to the same partite set of $G$.
Definition 3.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph and let A be the adjacency matrix of $G$. A column (say, the $j$-th column) is said to have almost consecutive ones if, between any two ones in that column, whenever a zero appears at the position $(v_{l},v_{j})$, the vertices $v_{l}$ and $v_{j}$ belong to the same partite set of $G$.
We now introduce another type of vertex ordering for $r$-partite graphs, referred to as the $r$-interval ordering. Using this ordering, we characterize the class of interval $r$-graphs.
Let $G=(X_{1},X_{2},\dots,X_{r},E)$ be an $r$-partite graph of order $n$, and let $v_{1},v_{2},\dots,v_{n}$ be an ordering of its vertices. Place these vertices on the real line such that the $i$-th vertex $v_{i}$ is placed on the positive integer point $i$.
Let A denote the adjacency matrix of $G$, where both rows and columns are arranged in the increasing order of vertex indices. Consider any row of A (say, the $i$-th row). Define $R_{i}$ as the set of entries of $1$ in this row that appear almost consecutively, beginning with the column corresponding to the vertex $v_{s_{i}}$, where $v_{s_{i}}$ is the first vertex to the right of $v_{i}$ that belongs to a different partite set and $s_{i}>i$. If no $v_{i}v_{s_{i}}\in E$, then $R_{i}=\emptyset$. Otherwise, the sequence continues rightward until the last $1$ in this almost consecutive manner is reached.
Similarly, consider any column of A (say, the $j$-th column). Define $C_{j}$ as the set of entries of $1$ in this column that appear almost consecutively, beginning with the row corresponding to the vertex $v_{s_{j}}$, where $v_{s_{j}}$ is the first vertex to the right of $v_{j}$ that belongs to a different partite set and $s_{j}>j$. If no $v_{j}v_{s_{j}}\in E$, then $C_{j}=\emptyset$. Otherwise, the sequence continues downward until the last $1$ in this almost consecutive manner is reached.
An ordering $v_{1},v_{2},\dots,v_{n}$ of the vertices of $G$ is called an $r$-interval ordering if the sets $R_{i}$ and $C_{j}$ together contain all the $1$’s in the adjacency matrix A.
Theorem 2.
An $r$-partite graph $G=(X_{1},X_{2},\dots,X_{r},E)$ is an interval $r$-graph if and only if its vertex set $X=\bigcup_{i=1}^{r}X_{i}$ admits an $r$-interval ordering.
Proof.
Necessity: Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an interval $r$-graph of order $n$. Then there exists an interval model
$\mathcal{I}=\{I_{v}:v\in X=\bigcup_{i=1}^{r}X_{i}\}$
such that for any two vertices $u,v$ belonging to different partite sets,
$uv\in E\quad\text{if and only if}\quad I_{u}\cap I_{v}\neq\emptyset.$
Without loss of generality, we may assume that the interval model $\mathcal{I}$ satisfies the following conditions:
  1. 1.
    All intervals are closed, i.e., each interval includes its endpoints.
  2. 2.
    No two intervals share the same left endpoint.
Label the vertices of $G$ as $v_{1},v_{2},\ldots,v_{n}$ according to the increasing order of the left endpoints of their corresponding intervals. Arrange both the rows and columns of the adjacency matrix $A$ of $G$ according to this vertex ordering.
We claim that, under this arrangement, the sets $R_{i}$ and the sets $C_{j}$ together contain all the 1’s of the adjacency matrix $A$.
Suppose an entry of $A$ is $a_{ij}=1$. This implies that $v_{i}$ and $v_{j}$ are adjacent in $G$, i.e., $v_{i}v_{j}\in E$. Let $v_{i}\in X_{\alpha}$ and $v_{j}\in X_{\beta}$, where $\alpha\neq\beta$. We distinguish two cases according to the relative positions of $v_{i}$ and $v_{j}$ in the vertex ordering.
Case 1:
When $i>j$. From the interval representation (see Figure 2), any vertex appearing between $v_{j}$ and $v_{i}$ in this ordering, and belonging to a partite set other than $X_{\beta}$, will be adjacent to $v_{j}$. Consequently, $C_{j}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
$v_{1}$$v_{2}$$v_{j}$$v_{j+1}$$v_{j+2}$$v_{i-1}$$v_{i}$$v_{i+1}$$v_{n-1}$$v_{n}$$I_{v_{i}}$$I_{v_{j}}$
Figure 2: $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where $i>j$.
Case 2:
When $i<j$. Similarly, any vertex appearing between $v_{i}$ and $v_{j}$ that does not belong to the same partite set $X_{\alpha}$ will be adjacent to $v_{i}$ (see Figure 3). Hence, $R_{i}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
$v_{1}$$v_{2}$$v_{i}$$v_{i+1}$$v_{i+2}$$v_{j-1}$$v_{j}$$v_{j+1}$$v_{n-1}$$v_{n}$$I_{v_{j}}$$I_{v_{i}}$
Figure 3: $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where $i<j$.
In either case, the 1 at the $(i,j)$-th entry is contained in one of $R_{i}$ or $C_{j}$. Therefore, the sets $R_{i}$ and $C_{j}$ together include all the 1’s of the adjacency matrix $A$.
Sufficiency: Consider an $r$-partite graph $G=(X_{1},X_{2},\ldots,X_{r},E)$, where the vertex set $X=\bigcup_{i=1}^{r}X_{i}$ is ordered as $v_{1},v_{2},\ldots,v_{n}$. Assume that this ordering is an $r$-interval ordering.
Assume if the rows and columns of the adjacency matrix of $G$ are arranged according to this ordering, then the sets $R_{i}$ and $C_{j}$ together contain all the 1’s in the adjacency matrix.
We now construct an interval model for $G$. If $R_{i}=\emptyset$, define $I_{v_{i}}=[i,i]$. Otherwise, let $r_{i}$ be the greatest column index such that the entry at the $(i,r_{i})$-th position of $A$ is 1. Then define $I_{v_{i}}=[i,r_{i}]$.
It remains to show that the family of intervals $\mathcal{I}=\{I_{v_{i}}:1\leq i\leq n\}$ serves as an interval model for the $r$-partite graph $G$.
Suppose $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where ($i<j$) and $v_{i}$, $v_{j}$ belong to different partite sets. Then by the construction of the intervals we can say that $j\in I_{v_{i}}$ $\implies\mathcal{R}_{i}$ contains the $1$ at position $(i,j)$ of the adjacency matrix of $G$. Therefore $v_{i}$ is adjacent to $v_{j}$.
Conversely, let $v_{i}$ is adjacent to $v_{j}$ in $G$, where $(i>j)$ , say. Therefore $\mathcal{R}_{j}$ will contain the 1 at the $(j,i)$-th entry. Implies $I_{v_{j}}$ contains $i$, hence by the definition of $I_{v_{i}}$ we have $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$. Therefore the class of intervals as constructed above is an interval model for the $r$-partite graph and so $G=(X_{1},X_{2},...,X_{r},E)$ is an interval $r$-graph.
Example 1. Consider the following $3$-partite graph in Figure 4. Using the above theorem we show that it is an interval $3$-graph.
$v_{7}$$v_{5}$$v_{3}$$v_{4}$$v_{6}$$v_{10}$$v_{1}$$v_{2}$$v_{9}$$v_{8}$
Figure 4: A 33-partite graph.
[…]
Figure 5: The adjacency matrix of the 3-partite graph in Figure 4, where the rows and columns are arranged according to the increasing order of their indices, and corresponding $\mathcal{R}_{i}$’s and $\mathcal{C}_{j}$’s are shown in the elliptic regions. Its interval representation is as follows: Red: $I_{v_{1}}=[1,3]$, $I_{v_{6}}=[6,9]$, $I_{v_{10}}=[10,10]$; Green: $I_{v_{2}}=[3,4]$, $I_{v_{8}}=[8,10]$, $I_{v_{9}}=[9,9]$; Black: $I_{v_{3}}=[3,3]$, $I_{v_{4}}=[4,6]$, $I_{v_{5}}=[5,8]$, $I_{v_{7}}=[7,9]$.
Note that the zeros inside the elliptic regions of $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$, as shown in Figure 5, do not belong to the sets $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$. They simply indicate that the 1’s in $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$ are not strictly consecutive, but rather almost consecutive.
Pavol Hell and Jing Huang [6] characterized interval bigraphs using forbidden patterns with respect to a specific vertex ordering in the following theorem.
Theorem 3 ([6]).
Let $H$ be a bipartite graph with bipartition $(X,Y)$. Then the following statements are equivalent:
  • $H$ is an interval bigraph;
  • the vertices of $H$ can be ordered $v_{1}$, $v_{2}$, …, $v_{n}$, so that there do not exist $a<b<c$ in the configurations in Figure 6. (Black vertices are in $X$, red vertices in $Y$, or conversely, and all edges not shown are absent.)
  • the vertices of $H$ can be ordered $v_{1}$, $v_{2}$, …, $v_{n}$, so that there do not exist $a<b<c<d$ in the configurations in Figure 7.
$v_{a}$$v_{b}$$v_{c}$
Figure 6: Forbidden pattern.
[…]
Figure 7: Forbidden patterns.
It is a natural question whether the classes of interval $r$-graphs ($r\geq 3$) can be characterized by a finite collection of forbidden patterns with respect to some specific ordering of their vertices. The following theorem provides an answer to this question.
Theorem 4.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph ($r\geq 3$). Then the following statements are equivalent:
  1. i.
    $G$ is an interval $r$-graph;
  2. ii.
    The vertices of $G$ can be ordered $v_{1},v_{2},\ldots,v_{n}$ such that no indices $i<j<k$ occur in the configurations shown in Figure 8. (Here, different colors indicate that vertices belong to different partite sets; moreover, all edges not explicitly drawn are absent.)
[…]
Figure 8: Forbidden patterns for interval $r$-graphs ($r\geq 3$) in terms of specific orderings of their vertices.
Proof.
Necessity. Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ $(r\geq 3)$ be an interval $r$-graph. Let $\mathcal{I}=\{I_{v}:v\in\bigcup_{i=1}^{r}X_{i}\}$ denote an interval model of $G$.
Without loss of generality, we may assume that the interval model $\mathcal{I}$ satisfies the following conditions:
  1. 1.
    All intervals are closed, i.e., each interval includes its endpoints.
  2. 2.
    No two intervals share the same left endpoint.
Arrange the vertices of $G$ according to the increasing order of the left endpoints of the corresponding intervals.
Consider any three vertices $v_{i},v_{j},v_{k}$ of $G$ belonging to three distinct partite sets $X_{\alpha},X_{\beta},X_{\gamma}$, respectively, where $i<j<k$. Suppose that $v_{i}$ and $v_{k}$ are adjacent. According to the vertex ordering, the left endpoint of the interval $I_{v_{k}}$ must lie within the interval $I_{v_{i}}$. Consequently, the left end points of the intervals $I_{v_{i+1}},I_{v_{i+2}},\ldots,I_{v_{k-1}}$ are contained in $I_{v_{i}}$, and therefore each of them intersects $I_{v_{i}}$.
It follows that every vertex indexed between $v_{i}$ and $v_{k}$ that does not belong to $X_{\alpha}$ must be adjacent to $v_{i}$. In particular, $v_{j}\in X_{\beta}$ is adjacent to $v_{i}$. Hence, the configurations (ii) and (iii) shown in Figure 8, cannot occur in an interval $r$-graph.
Similarly, if the vertices $v_{j}$ and $v_{k}$ belong to the same partite set, then the configuration (i) illustrated in Figure 8, also cannot occur in an interval $r$-graph. Thus, none of the three configurations shown in Figure 8 can occur in $G$.
Sufficiency: Suppose $G=(X_{1},X_{2},\ldots,X_{r},E)$ is an $r$-partite graph that admits an ordering $v_{1},v_{2},\ldots,v_{n}$ such that none of the configurations illustrated in Figure 8 occur. We shall show that $G$ is an interval $r$-graph.
Place the vertices of $G$ on the real line such that each vertex $v_{i}$ is positioned at the positive integer point $i$, for $1\leq i\leq n$. We now construct an interval corresponding to each vertex $v_{i}$ ($1\leq i\leq n$).
For a vertex $v_{i}\in X_{\alpha}$, define $t_{i}$ to be the largest vertex index (in the rightward direction from $v_{i}$) corresponding to a vertex $v_{t_{i}}\in X\setminus X_{\alpha}$ such that all vertices $v_{l}\in X\setminus X_{\alpha}$, where $i<l\leq t_{i}$, are adjacent to $v_{i}$. If no such vertex index exists, then set $t_{i}=i$.
Now define a family of intervals
$\mathcal{I}^{\prime}=\{I^{\prime}_{v_{i}}:1\leq i\leq n\}$
corresponding to the vertex set of $G$, where
$I^{\prime}_{v_{i}}=[i,\,t_{i}+\tfrac{1}{2}].$
We will show that this family of intervals is an interval representation of the graph $G$.
Claim 1. If two vertices $v_{i}\in X_{\alpha}$ and $v_{k}\in X_{\gamma}$ (with $\alpha\neq\gamma$ and $i<k$) satisfy $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$, then $v_{i}$ and $v_{k}$ are adjacent in $G$.
Proof. From the definition of the intervals,
$I^{\prime}_{v_{i}}=[i,t_{i}+\tfrac{1}{2}],\quad I^{\prime}_{v_{k}}=[k,t_{k}+\tfrac{1}{2}].$
If $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$, then $k\in I^{\prime}_{v_{i}}$, implying
$i\leq k\leq t_{i}+\tfrac{1}{2}\implies i\leq k\leq t_{i}.$
Hence, $v_{i}$ is adjacent to $v_{k}$.
Claim 2. Conversely, suppose $v_{i}$ and $v_{k}$ ($i<k$) are adjacent in $G$, where $v_{i}\in X_{\alpha}$ and $v_{k}\in X_{\gamma}$. We will show that $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$.
Proof. For contradiction, assume that $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}=\emptyset$. Then, by definition of the intervals, $t_{i}<k$. This means that $v_{i}$ is not adjacent to every vertex consecutively (from partite sets different from $X_{\alpha}$) on its right up to $v_{k}$. Hence, there exists a vertex $v_{j}$ such that $v_{j}\notin X_{\alpha}$, $i<j<k$, and $v_{i}$ is not adjacent to $v_{j}$.
We consider the following cases:
  • Case 1. $v_{j}$ and $v_{k}$ belong to the same partite set $X_{\gamma}$. Then configuration (i) (see Figure 8) occurs — a contradiction.
  • Case 2. $v_{i},v_{j},v_{k}$ belong to three distinct partite sets $X_{\alpha},X_{\beta},X_{\gamma}$ respectively.
    • Subcase 1: If $v_{j}$ is not adjacent to $v_{k}$, then configuration (ii) (see Figure 8) occurs — a contradiction.
    • Subcase 2: If $v_{j}$ is adjacent to $v_{k}$, then configuration (iii) (see Figure 8) occurs — a contradiction.
Thus, in every possible case, we reach a contradiction. Therefore, if $v_{i}$ and $v_{k}$ are adjacent, then $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$.
Hence, the constructed family of intervals $\mathcal{I}^{\prime}$ serves as an interval model of the $r$-partite graph $G$. Therefore, $G$ is an interval $r$-graph.
3 Conclusion
In this paper, we discussed various types of vertex orderings for the most generalized version of interval bigraphs, namely interval $r$-graphs. We also identified the corresponding forbidden patterns for this class with respect to certain specific orderings. For future research, one could aim to extend these concepts to the class of circular-arc graphs, with the goal of generalizing the notion of circular-arc bigraphs as well. It would also be interesting to investigate whether a finite set of forbidden patterns exists for these generalized classes.
Acknowledgement
The first author sincerely acknowledges the Council of Scientific and Industrial Research (CSIR), India [File No. 09/0028(11986)/2021-EMR-I], for providing financial support through a fellowship to pursue the Ph.D. degree.